OpenStudy (anonymous):

find the exact value, in radians, of the expression tan-1(-square root 3)

OpenStudy (anonymous):

\[\text{ArcTan}\left[-\sqrt{3}\right]=-\frac{\pi }{3} \]

OpenStudy (anonymous):

how did you get to that answer?

OpenStudy (anonymous):

It is the special right triangle where the acute angles are 30 and 60 degrees. pi/3 is 60 degrees.

OpenStudy (anonymous):

wow thank you soo much

OpenStudy (anonymous):



OpenStudy (anonymous):

@Kiara_Wright10 Sorry. My bad. The idea was correct, the drawing was incorrect. Refer to the revised drawing above.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome.